a baseball is thrown straight upward with a speed of 30 m/s. (a) How long will it rise? (b) How high will it rise? (c) How long after it leaves the hand will it return to the starting point? (d) When will its speed be 16 m/s?


1
Expert's answer
2022-02-22T10:30:20-0500

Given:

"v_0=30\;\rm m/s"

"g=9.8\:\rm m/s^2"


(a) How long will it rise?

"t=\frac{v_0}{g}=\frac{30}{9.8}=3.1\;\rm s"


(b) How high will it rise?


"h_{\max}=\frac{v_0^2}{2g}=\frac{30^2}{2*9.8}=46\;\rm m"

(c) How long after it leaves the hand will it return to the starting point?

"t_1=2t=2*6.2\;\rm s"

(d) When will its speed be 16 m/s?

"t=\frac{v-v_0}{-g}=\frac{16-30}{-9.8}=1.4\:\rm s"